OpenStudy (narissa):

6(a+3)=18+a i think its no solution.

OpenStudy (narissa):

@texaschic101

OpenStudy (nykoreyan):

a=0

OpenStudy (narissa):

how did u get that answer? :)

OpenStudy (nykoreyan):

Solve for a by simplifying both sides of the equation, then isolating the variable

OpenStudy (texaschic101):

6(a + 3) = 18 + a --- distribute through the parenthesis 6a + 18 = 18 + a -- subtract a from both sides 6a - a + 18 = 18 -- subtract 18 from both sides 5a = 18 - 18 5a = 0 a = 0/5 or just 0

OpenStudy (nykoreyan):

which is what I said

OpenStudy (narissa):

there isnt at 0 as my answer choice

OpenStudy (texaschic101):

what are your answer choices ?

OpenStudy (texaschic101):

that does not make sense because if you sub in 0 for a, the equation is correct. 6(0 + 3) = 18 + 0 6(3) = 18 18 = 18 (correct)

OpenStudy (narissa):

A-1 B -3 C no soluion D all real numbers

OpenStudy (nykoreyan):

c no solution ............o and no solution is the same thing lol @narissa

OpenStudy (texaschic101):

Is your problem written correctly ?

OpenStudy (narissa):

oohh okay lol sorry i was a bit confused @nykoreyan

OpenStudy (babylove18):

its no solution you got it narissa

OpenStudy (texaschic101):

0 and no solution do not mean the same thing

OpenStudy (texaschic101):

but I guess I would put no solution...if I had to pick one

OpenStudy (nykoreyan):

well in this situation they would be the same thing